# Diffusion-Reinforcement Learning Hierarchical Motion Planning in Adversarial Multi-agent Games

### v1

@@ VI Conclusion and Future Work
In conclusion, we have presented a novel learning-based framework for evaders in large, partially observable, multi-agent pursuit-evasion settings. Our approach can help the evader reach the hideout and best conceal itself from capturing by using a hierarchy with a high-level diffusion path planner and a low-level RL evasive policy. Our work can also be extended to more general navigation problems (e.g. visual navigation, manipulator grasping etc.). However, our learned hierarchy requires additional interactions with the environments when constructing the costmap. In the future, we plan to integrate the map construction process into the RL training loop by iterating a location-only based value function $V(s)$ with the Bellman equation. Additionally, we will train the RL policy for both the pursuit and evasion sides simultaneously and investigate the relationship between the Nash equilibrium with respect to team composition.

### v2

@@ VI Conclusion and Future Work
In conclusion, we have presented a novel learning-based framework for evaders in large, partially observable, multi-agent pursuit-evasion settings. Our approach can help the evader reach the hideout and best conceal itself from capturing by using a hierarchy with a high-level diffusion path planner and a low-level RL evasive policy. In the future, we plan to train the whole framework end-to-end and bring our diffusion-RL technique to real-world navigation tasks involving high-level human domain knowledge transfer and low-level dynamic constraints (e.g. environment-aware path generation and finetuning in autonomous driving).